# Deep in the Knotted Black Hole

### v1

@@ I Introduction
Recently there has been growing interest in studying the response of detectors that freely fall into a black hole and cross its horizon. Notwithstanding the general expectation (rooted in the equivalence principle) that such a detector will have a smooth rate of response and will not thermalize, little is known about this case. A detector freely falling through a stationary cavity in a ($3+1$)-dimensional Schwarzschild background was found to have a different response from that of an equivalently accelerated detector travelling through an inertial cavity in flat spacetime [16]. The first numerical study of a detector freely falling toward and across the horizon of a Schwarzschild black hole [17] found – for both the Hartle-Hawking(-Israel) and Unruh states – that the response is smooth, and that it is monotonically increasing along the trajectory, except for a small but discernible local extremum near the horizon-crossing. There was however no analytic corroboration for the extremum, suggesting a need for further numerical checks [18].
@@ II.4 The Wightman function for the AdS, BTZ and geon spacetimes
We consider a detector interacting with a conformally coupled massless scalar field $\phi(x)$, that is, the field satisfies the Klein-Gordon equation $(\Box-R/8)\phi(x)=0$, where $\Box$ is the d’Alembert operator and $R=-6/\ell^{2}$ is the Ricci scalar.
@@
On the BTZ spacetime, $W_{\text{BTZ}}$ has the characteristics of a Hartle-Hawking-Israel (HHI) vacuum [37, 38], describing a black hole in thermal equilibrium with a heat bath at infinity [35]. On the geon spacetime, $W_{\text{geon}}$ is not a thermal equilibrium state as it is not invariant under the exterior Killing time translations, but it still has some characteristics of a HHI state [7, 11, 12, 39].
When the points $x$ and $x^{\prime}$ are in the exterior region, equations (8), (9), (14) and (16) give
@@
where $\Delta\phi=\phi-\phi^{\prime}$ and $\Delta t=t-t^{\prime}$, and the $\epsilon\to 0^{+}$ limit in (18a) specifies the distributional character of the Wightman function. Note that (18a) depends on $t$ and $t^{\prime}$ only via the combination $\Delta t$, showing explicitly that $W_{\text{BTZ}}$ is invariant under the exterior Killing time translations, but (18b) depends on $t$ and $t^{\prime}$ only via the combination $t+t^{\prime}$, showing that $W_{\text{geon}}$ is not invariant under the exterior Killing time translations.
@@ Acknowledgements
We thank Chris Shallue and Sean Carroll for sharing with us an early copy of their work on the Schwarzschild infall [18]. This work was supported in part by the Natural Sciences and Engineering Research Council of Canada. MRPR gratefully acknowledges the support provided by the Mike and Ophelia Lazaridis Graduate Fellowship. The work of JL was supported by United Kingdom Research and Innovation Science and Technology Facilities Council [grant numbers ST/S002227/1, ST/T006900/1 and ST/Y004523/1]. For the purpose of open access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript version arising.

### v2

@@ I Introduction
There has been growing interest in the response of detectors that freely fall into a black hole and cross its horizon. There appears to be a general expectation (perhaps rooted in the equivalence principle) that such a detector will not thermalize, but characterising the detector’s response more precisely is a topic of continuing study. An adiabatic formalism for estimating the detector’s effective temperature has been developed in [16, 17, 18, 19, 20]. A detector freely falling through a stationary cavity in a ($3+1$)-dimensional Schwarzschild background was found to have a different response from that of an equivalently accelerated detector travelling through an inertial cavity in flat spacetime [21]. A detector falling freely toward and across the horizon of a Schwarzschild black hole has been analysed numerically in [22, 23]. An operational definition for an effective temperature that accommodates the brief time that the detector spends near the horizon was proposed in [23].
@@ II.4 The Wightman function for the AdS, BTZ and geon spacetimes
We consider a detector interacting with a conformally coupled massless scalar field $\phi(x)$, that is, the field satisfies the Klein-Gordon equation $(\Box-R/8)\phi(x)=0$, where $\Box$ is the d’Alembert operator and $R=-6/\ell^{2}$ is the Ricci scalar [40]. The numerical coefficient $1/8$ is the conformal coupling value $\frac{1}{4}(d-2)/(d-1)$ in $d=3$ spacetime dimensions.
@@
$W_{\text{BTZ}}$ has the characteristics of a Hartle-Hawking-Israel (HHI) vacuum [43, 44], describing a black hole in thermal equilibrium with a heat bath at infinity [41]. In particular, the restriction of $W_{\text{BTZ}}$ to an exterior region of the BTZ spacetime is stationary, because the map $\Gamma$ (9) on the AdS3-Rindler spacetime commutes with the Killing vector $r_{h}\ell^{-2}(V\partial_{V}-U\partial_{U})$ that generates the exterior Killing time translations. By contrast, $W_{\text{geon}}$ does not describe a thermal equilibrium state on the geon. The second term in (17b) shows that $W_{\text{geon}}$ differs from $W_{\text{BTZ}}$ already in the geon’s exterior, and this term further shows that the restriction of $W_{\text{geon}}$ to the geon’s exterior is not invariant under the exterior Killing time translations, since the quotienting map $J$ (14) does not commute with the BTZ Killing time translations. It hence follows that $W_{\text{geon}}$ cannot describe a thermal equilibrium state, although it still has some characteristics of a HHI state, as analysed in [7, 11, 12, 45, 8]. In particular, the response of a static detector in the geon’s exterior is not stationary, in contrast to the response of a static detector in the BTZ exterior, although the two asymptotically agree for a detector that operates at times much earlier than the $t=0$ hypersurface shown in Figure 2 [8].
We shall proceed to investigate how the responses differ for an infalling detector. As preparation, we record here that when the points $x$ and $x^{\prime}$ are in the exterior region, equations (8), (9), (14) and (16) give
@@
where $\Delta\phi=\phi-\phi^{\prime}$ and $\Delta t=t-t^{\prime}$, and the $\epsilon\to 0^{+}$ limit in (18a) specifies the distributional character of the Wightman function. Note that (18a) depends on $t$ and $t^{\prime}$ only via the combination $\Delta t$, whereas (18b) depends on $t$ and $t^{\prime}$ only via the combination $t+t^{\prime}$. This shows explicitly that $W_{\text{BTZ}}$ is invariant under the exterior Killing time translations but $W_{\text{geon}}$ is not.
@@ Acknowledgements
We thank Chris Shallue and Sean Carroll for sharing with us an early copy of their work on the Schwarzschild infall [23]. We thank an anonymous referee for helpful presentational suggestions. This work was supported in part by the Natural Sciences and Engineering Research Council of Canada. MRPR gratefully acknowledges the support provided by the Mike and Ophelia Lazaridis Graduate Fellowship. The work of JL was supported by United Kingdom Research and Innovation Science and Technology Facilities Council [grant numbers ST/S002227/1, ST/T006900/1 and ST/Y004523/1]. For the purpose of open access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript version arising.